# The Great Comets of 1843 and 1882 at Their Previous Return to Perihelion in the Twelfth Century: One Spectacular, the Other Dull

### v1

@@ 6. Comet X/1106 C1 at and After Sunset
Although the head of the comet was not bright enough to be detected at sunset, Figure 5 shows that the observing conditions at the northern latitudes near 30 were rapidly improving with time in early February. This applies not only to the Chinese observations, but also to those from Palestine (e.g., the city of Jerusalem is at nearly the same latitude) and other locations. The large recorded disparities in the tail length could be explained by the wide variety in both the sites and times of observation. Reports of shorter tail lengths could be expected from more northerly locations and from later, post-twilight observations, when the comet’s head was deep below the horizon and only distant parts of the tail could still be seen. This means that the best estimates of the tail length are the longest ones, near 100. The reports that the comet was “split into many pieces,” or “appeared like a broken-up star,” or like “a beam forking into many rays,” provide likely evidence for the presence of a plasma tail and its complex morphology. Although in the equatorial coordinate system the predicted position of the plasma tail was in the east-southeast direction (varying from position angle of 116 on February 5 to 103 on February 13), in the horizontal system at latitude ${}^{\>\!\!\circ}$ the tail should have been seen to point up and east, deviating at sunset from the zenith direction by 63 on February 5 and by 55 on February 13.
@@
To depict the observing conditions as the 1106 comet was approaching the horizon following the sunset, I plot in Figure 6 its apparent motion over a period of 80 minutes on the evening of February 9, predicted for a location at geographic longitude of 120E and latitude of ${}^{\>\!\!\circ}$. At the sunset, which was found to have taken place at 9:45 UT, the comet was ${}^{\circ}\!$.1 from the Sun at position angle of 108 and elevation of ${}^{\circ}\!$.4. The comet set 67 minutes later, in the course of astronomical twilight, with the Sun 14 below the horizon. Under these conditions the long tail must have offered a magnificent spectacle.
The potential visibility of the comet’s head in twilight was another issue of interest. I already noted that the head was too faint to detect at sunset. As the Sun was moving deeper below the horizon, the conditions were improving. Unfortunately, the comet’s elevation was also diminishing. One would expect that the optimum conditions may take place somewhere midway between sunset and setting of the comet, a guess that is supported by quantitative examination in Table 4. For the perihelion time of 1106 February 2.25 and the selected observing site in China, the post-sunset conditions were compared for February 9 (sunset at 9:45 UT) and February 13 (sunset at 9:48 UT). The predicted apparent magnitudes of the comet were $-2.5$ mag and $-0.8$ mag, respectively, and the effects of interference by the Moon were accounted for; it was 3.5 days old and 27 from the comet’s head on February 9, 7.5 days old and 68 away on the 13th. The table shows that the comet’s head is predicted to have been visible to the naked eye over a period of 40 minutes (from 15 to 55 minutes after sunset) on February 9 and over a period of 77 minutes (from 17 to 94 minutes after sunset) on February 13. At the most favorable times on the two days the comet’s head was found to be, respectively, a little over 2 mags and 3 mags above the naked-eye visibility threshold.
@@ 7. The Comet of 1138 at and before Sunrise
Of distinct assistance to the projected dust tail’s improved visibility was the Earth’s transit across the comet’s orbit plane on August 20, which helped increase its surface brightness; this was the intervening event mentioned above under (ii). As a result, the line of sight from the Earth made only ${}^{\circ}\!$.7 with the orbit plane at the time of observation. Also, because of the orbit plane crossing, the tail was curving in the opposite sense on September 2 (Figure 7) compared to August 7 (Figure 4).
@@
Because all dust ejected about 0.1 day after perihelion and later is seen in Figure 7 to have by early September been projecting at larger distances from the Sun and at greater elevations than the comet’s head, it was this part of the tail that had the best chance of detection at that time. This is my argument under point (iii) above. In particular, one can see from the figure that the dust ejected near 1 day after perihelion was in fact located at an elevation about 6 higher in the sky than the head. It is hard to estimate how close to the horizon the feature could have been followed, but chances were that the visible tail length was at most 10 or so. It is also likely that the predicted marginal naked-eye detection of the head (Table 6) meant in practice that it did not stand out enough to be readily noticed. Its position, sideways from the main body of the dust tail and near the middle along its long axis, was very unusual and to the medieval observer could have been puzzling to the point that he probably would have commented on it.
The most unusual conditions of tail projection in September are further underlined by the data presented in Table 7, which likewise assist in estimating the tail length in early September and also help explain why the comet was seen over a rather limited period of time. The striking features of the table are the enormous distances from the Sun and even more so from the Earth reached by the submicron-sized dust particles (subjected to a radiation pressure acceleration of 0.6 the solar gravitation acceleration) ejected in the first hours after perihelion. (Such grains dominated the tails of the great Kreutz sungrazers, such as Ikeya-Seki, early after perihelion.) For example, the particles ejected a little less than one hour post-perihelion, which in Figure 7 make up the part of the tail appearing just barely above the horizon, were located more than 6 AU from the Sun and more than 7 AU from the Earth on September 2 UT and between 11 and 12 AU on September 28 UT, i.e., the morning hours of September 29 local time, when the comet went out of sight (Ho 1962). On the other hand, the heliocentric and geocentric distances of the microscopic dust ejected as late as 24 hours after perihelion, which occupies the area of the tail near its highest point in Figure 7, were only between 2 and 2.7 AU on September 2 and between 3.6 and 3.9 AU on September 28; of course, the production rate of dust was much higher one hour after perihelion than 23 hours later.
@@
In this context, comparison with the sungrazing comet Pereyra (C/1963 R1) is illuminating. It passed perihelion on September 23, merely three weeks later in the year than predicted for the comet of 1138. Accordingly, Pereyra’s observing conditions were more favorable, but not by much. Also, Pereyra was not a member of Population II, so that the geometry was slightly different. Yet, while intrinsically brighter than Ikeya-Seki, [^6] From sets of visual brightness observations I determined the post-perihelion absolute magnitudes of 5.9 for Ikeya-Seki and 4.4 for Pereyra (Sekanina 2002); these compare with an adopted absolute magnitude of 0.5 for the comet of 1138. Pereyra was not detected in daylight, [^7] If the arrival of comet Pereyra were delayed by 16 years (less than 2 percent of its orbital period), it would have saturated the imager on board the Solwind satellite. but was discovered three weeks after perihelion. Over a period of several days following its discovery, Pereyra was a naked-eye object, displaying a tail several degrees in length. Although the conditions were clearly more favorable in the southern hemisphere (e.g., Bennett et al. 1964), the comet and its tail were seen by northern-hemisphere observers as well. [^8] The visibility of comet Pereyra with the naked eye is apparent from the narratives by some observers [e.g., Capen (1964), who was stationed at Wrightwood, California], but also from an incredible story involving a 1977 letter written by G. E. D. Alcock, in which — as aptly recounted in 2003 by B. G.,Marsden in his lecture given in Alcock’s memory to a RAS-BAA Pro-Am Discussion Meeting: Meteorites, Meteors, & Comets and documented by Shanklin (2004) — he wondered, in retrospect, whether a thin pencil-like beam of light that he had seen low in the sky on 1963 September 12, two days before comet Pereyra was discovered, happened to be the comet’s tail. Telescopically the comet was under observation over a period of three months, with the last observation in mid-December made from the northern hemisphere. Thus, notwithstanding the various differences (such as orbital incompatibility, a gap of three weeks in the perihelion time, an intrinsic-brightness gap of almost 4 magnitudes, etc.), one notices very obvious similarities between the predicted traits of the 1138 comet and the observed traits of comet Pereyra.
@@
The second conclusion concerns a rule of thumb to investigate a correlation between the time of last naked-eye observation and the comet’s apparent magnitude. Sightings of a bright sungrazing comet in the ancient or medieval times were those of its tail, which was seen even after the comet’s head was no longer visible to the unaided eye. I tested this relationship on the well-observed sungrazers, such as the Great March Comet of 1843, the Great September Comet of 1882, Ikeya-Seki, etc., and suggested that, based on the experience with these objects, the post-perihelion tail of a sungrazer was typically seen with the naked eye until the brightness faded to apparent magnitude of $\sim$7, so that the head became a telescopic object long before the last naked-eye tail detection (Sekanina 2022).
@@
Applied to seven objects between AD 943 and 1695, listed by Hasegawa & Nakano (2001) as potential Kreutz sungrazers (all reaching perihelion under favorable observing conditions in Octover or November, similar to those for comet Ikeya-Seki in 1965), this exercise led to a revision of their absolute brightness downward, the correction averaging more than 2 magnitudes. The outcome was their better agreement with the absolute magnitudes of the recent Kreutz sungrazers.
@@
A very different case on that same list is the comet of 1041, which was already briefly commented on in Section 2. Hasegawa & Nakano assigned it an absolute magnitude of --1, making it 2 mags intrinsically brighter than was my estimate for the Kreutz system’s progenitor (Sekanina 2022)! There appear to be two problems with this 1041 object, the more serious one being that two different comets — either one recorded in China as well as Korea — were almost certainly involved, as suggested by both Ho (1962) and England (2002). The first of the two comets was discovered on September 1 (Hasegawa & Nakano 2001) and “went out of sight after more than 20 days” (Ho 1962). The mentioned rule-of-thumb method offered for it absolute magnitude 4, which is 5 mags fainter than proposed by Hasegawa & Nakano. The second comet appeared in November, was seen for more than 10 days, and appears to have been intrinsically fainter. Both may have been Kreutz sungrazers, perhaps even fragments of an object that split not long before 1041.
@@
The estimated (post-perihelion) absolute magnitude 4 for the September comet of 1041 happens to be brighter than Pereyra’s by only 0.5 mag or less and too faint for the parent, especially because circumstantial evidence suggests that the 1041 comet fragmented heavily at perihelion. [^9] The barycentric orbital period of the 1041 comet upon its arrival at perihelion was 678 years, but Pereyra’s original orbital period slightly exceeded 900 years, which implied the existence of a few fragments in an orbital-period range of $\sim$450–750 years. This would reduce the comet’s rate of post-perihelion fading, thus bringing the relationship between the absolute magnitudes of the 1041 comet and Pereyra closer to expectation. If so, its brightness should have been almost comparable to the brightness of the 1106 or 1138 comets. To achieve this, the time of the last naked-eye observation in the rule of thumb should have coincided with the comet’s apparent magnitude having equaled not 7, but about 4.5. And this is the second problem with the September comet of 1041.
@@ 8. Fragmented Nucleus of the Comet of 1138
Even though I believe that the nucleus of the Great Comet of 1106 did break up [^10] It is possible that besides the Great March Comet of 1843, at least one other sungrazer — C/1668 E1 — was also a fragment of the 1106 comet (Sekanina 2025). in the course of its perihelion passage as well, only the fragmentation event experienced by the Chinese comet of 1138 is modeled and examined here in some detail. There have been two good reasons: (i) it is virtually certain that the Great September Comet of 1882 and Ikeya-Seki were products of this event (cf. Marsden 1967), possibly with other objects; and (ii) the nuclei of the 1882 comet and Ikeya-Seki were themselves subject to fragmentation at perihelion, the first of the two splitting into five to six components; a search for similarities has been desirable.
The multiple nuclei of the 1882 comet, as well as the existence of split comets whose breakup is known to be of tidal nature but occurred at close proximity of Jupiter — such as D/1993 F2 (Shoemaker-Levy) or 16P/Brooks (C/1889 N1) — suggest that (a) the formation of a chain of more than two nearly equidistant fragments, buried in an oval cloud of debris, is by no means exceptional; and (b) as a rule, the principal fragment is — unlike in nontidally split comets — located typically not at the leading end of the chain, but rather nearer its middle. The two conditions combine to become helpful for predictive purposes. In applications, the first condition serves as an impetus in a search for conceivable fragments among potential historical sungrazers, whereas — once the principal fragment has been identified (or probably identified) — the second condition is useful in assessing a timeline of the chain of fragments, with the aim to constrain the range of fragments’ orbital periods as much as possible. Of particular interest is any hint of evidence that additional fragments could be expected to arrive in the future, especially near future.
@@
In the context of the conditions (a) and (b) above, the adherence to these postulates led to the conclusion that at close proxinity to perihelion the Sun’s tidal force did, in all probability, split the nucleus of the 1138 comet not into two, but at least four or perhaps five or six components (Sekanina 2025). Chronologically, such a chain of fragments could have started with X/1702 D1 and continued with the “sun-comet” of 1792 (Strom 2002), [^12] Because of very vague nature of the “sun-comets,” this obviously is a weak entry; it is possible that the fragment expected to arrive at about this time was missed. followed by the 1882 sungrazer (the principal, most massive fragment) and by Ikeya-Seki. In order that the principal fragment be located near the middle of the chain, in line with condition (b), one can argue that at least one additional fragment should exist beyond Ikeya-Seki, and it should arrive in the near future. There is a good chance that this could happen — with the uncertainty of at least several years — around 2027, while the next one (if any) should be a matter of distant future.
The appearance of the nucleus region of a split sungrazer is essentially determined by the angular separations between the major, kilometer-sized and larger fragments, which are buried in an encompassing cloud of debris made up of smaller fragments down to microscopic dust. To describe the feature’s overall dimensions, one needs to establish the osculating orbits of the major fragments. Since a tidal breakup that involves no significant momentum exchange among the major fragments has practically no effect on their perihelion times and angular elements, they remain equal to those of the parent at the time of breakup. For the 1138 comet they are listed in Table 1. The remaining two elements vary from fragment to fragment, but the perihelion distances marginally, by a fraction of the nucleus dimensions. The only meaningful differences are in the fragments’ eccentricities, $e$, or alternatively in their semimajor axes, $a$.
The starting point in the process of modeling the appearance of the 1138 comet’s fragmented nucleus is the relationship between its osculating and future orbital periods as the previous appearance of the 1882 sungrazer, its principal fragment. From Table 1 the relevant osculating period is 754.4 yr, while the future barycentric orbital period is 1882.7 – 1138.6 = 744.1 yr. [^13] The actual calculation of the period has to be done in Julian dates (not in calendar days), which are then converted to years. This is necessary because the perihelion times in 1138 and 1882 are given, respectively, in the Julian and Gregorian calendars. When rounded off to a tenth of the year, the difference does not show up. In terms of $1/a$, the effect of the planetary perturbations of the 1882 sungrazer’s motion is
@@
Figure 8.— Modeled appearance of the fragmented nucleus of the 1138 comet on 1138 September 2.85 UT, nearly 33 days after perihelion. Plotted are the astrometric positions (equinox J2000) of the proposed major fragments, besides the Great September Comet of 1882 and Ikeya-Seki also the assumed fragments X/1702 D1, and the 1792 sun-comet; and a predicted sungrazer of 2027. The extent of the chain amounts to ${}^{\prime\prime}\!$.6 in position angle of ${}^{\circ}\!$.3 and is enveloped in an elongated cloud of dust debris. All five fragments are assumed to be active at the time to a degree that their condensations are not overwhelmed by the diffused light from the debris, an ideal case discussed below.
@@
The last remaining step is to compute the astrometric positions of the five fragments for the time of interest. Displayed in Figure 8 is the modeled appearance of the nucleus region of the 1138 comet on September 3 around 4:30 local time (September 2.85 UT). The extent of the fragmented nucleus (from X/1702 D1 to the predicted comet of 2027) amounts to ${}^{\prime\prime}\!$.6 in position angle of ${}^{\circ}\!$.3. Terminating the chain of fragments instead with Ikeya-Seki, the length would at the time be ${}^{\prime\prime}\!$.2, with Ikeya-Seki and the 1882 sungrazer only ${}^{\prime\prime}\!$.8 apart. At the geocentric distance of 1.76 AU this was equivalent to merely 3600 km, but because the line of sight made with the line connecting the two objects an angle of merely 14, they were in space 14,500 km apart.
It is illuminating to compare the modeled appearance of the 1138 comet in Figure 8, at a heliocentric distance of 1.12 AU, with the observed appearance of the fragmented nucleus of the 1882 sungrazer at similar post-perihelion distances. Since it will never be known how the nucleus of the 1138 comet actually looked like on September 2 UT, this exercise is purely academic. Yet, inspection of the range of possible appearances offers an insight into the potential subsequent evolution and ultimate fate of the fragments.
@@
The first of these reproductions, in Figure 9, shows the fragmented nucleus on 1882 October 13 UT, 25.4 days after perihelion, as seen by Elgin with a 15-cm Grubb equatoreal. The extent of about 20′′ exceeds by 5′′ the length of the modeled chain of fragments of the 1138 comet. As Table 8 explains, this apparent imbalance happens to be the effect partly of a larger geocentric distance of the 1138 object, partly of its highly unfavorable projection, as the chain of fragments made an angle of only 14 with the line of sight. In reality, the 1882 fragmented nucleus was 38,000 km long at the time, less than one half the modeled 1138 nucleus.

### v2

@@ 6. Comet X/1106 C1 at and After Sunset
Although the head of the comet was not bright enough to be detected at sunset, Figure 5 shows that the observing conditions at the northern latitudes near 30 were rapidly improving with time in early February. This applies not only to the Chinese observations, but also to those from Palestine (e.g., the city of Jerusalem is at nearly the same latitude) and other locations. The large recorded disparities in the tail length could be explained by the wide variety in both the sites and times of observation. Reports of shorter tail lengths could be expected from more northerly locations and from later, post-twilight observations, when the comet’s head was deep below the horizon and only distant parts of the tail could still be seen. This means that the best estimates of the tail length were the longest ones, near 100. The reports that the comet was “split into many pieces,” or “appeared like a broken-up star,” or like “a beam forking into many rays,” provide likely evidence for the presence of a plasma tail and its complex morphology. Although in the equatorial coordinate system the predicted position of the plasma tail was in the east-southeast direction (varying from position angle of 116 on February 5 to 103 on February 13), in the horizontal system at latitude ${}^{\>\!\!\circ}$ the tail should have been seen to point up and east, deviating at sunset from the zenith direction by 63 on February 5 and by 55 on February 13.
@@
To depict the observing conditions as the 1106 comet was approaching the horizon following the sunset, I plot in Figure 6 its apparent motion over a period of 80 minutes on the evening of February 9, predicted for a location at geographic longitude of 120E and latitude of ${}^{\>\!\!\circ}$. At sunset, which was calculated to have taken place at 9:45 UT, the comet was ${}^{\circ}\!$.1 from the Sun at position angle of 108 and elevation of ${}^{\circ}\!$.4. The comet set 67 minutes later, in the course of astronomical twilight, with the Sun 14 below the horizon. Under these conditions the long tail must have offered a magnificent spectacle.
The potential visibility of the comet’s head in twilight was another issue of interest. I already noted that the head was too faint to detect at sunset. As the Sun was moving deeper below the horizon, the conditions were improving. Unfortunately, the comet’s elevation was also diminishing. One would expect that optimum conditions may have taken place somewhere midway between sunset and setting of the comet, a guess that is supported by quantitative examination in Table 4. For the perihelion time of 1106 February 2.25 and the selected observing site in China, the post-sunset conditions were compared for February 9 (sunset at 9:45 UT) and February 13 (sunset at 9:48 UT). The predicted apparent magnitudes of the comet were $-3.0$ mag and $-1.3$ mag, respectively, and the effects of interference by the Moon were accounted for; it was 3.5 days old and 27 from the comet’s head on February 9, 7.5 days old and 68 away on the 13th. The table shows that the comet’s head is predicted to have been visible to the naked eye over a period of 45 minutes (from 12 to 57 minutes after sunset) on February 9 and over a period of 81 minutes (from 14 to 95 minutes after sunset) on February 13. At the most favorable times on the two days the comet’s head is found to have been, respectively, 2.8 mags and 3.5 mags above the naked-eye visibility threshold.
@@ 7. The Comet of 1138 at and before Sunrise
Of distinct assistance to the dust tail’s improved visibility was the Earth’s transit across the comet’s orbital plane on August 20, increasing the tail’s surface brightness; this was the intervening event under (ii). As a result, the line of sight from the Earth made only ${}^{\circ}\!$.7 with the orbital plane at the time of observation on September 3. Also, because of the orbital plane crossing, the tail was curving in the opposite sense in early September (Figure 7) compared to August 7 (Figure 4).
@@
Because all dust ejected later than about 0.1 day after perihelion is seen in Figure 7 to have by early September been projecting at larger distances from the Sun and at greater elevations than the comet’s head, it was this part of the tail that had the best chance of detection at that time. This is my argument under point (iii) above. In particular, one can see from the figure that the dust ejected near 1 day after perihelion was in fact located at an elevation about 6 higher in the sky than the head. It is hard to estimate how close to the horizon the feature could have been followed, but chances were that the visible tail length was at most 10 or so. It is also likely that the predicted marginal naked-eye detection of the head (Table 6) meant in practice that it did not stand out enough to be readily noticed. Its position, sideways from the main body of the dust tail and near the middle along its long axis, was very unusual and to the medieval observer puzzling to the point that, if seen, he probably would have commented on it.
The highly unusual conditions of tail projection are likewise underlined by the data presented in Table 7, which further assist in estimating the tail length in early September and also help explain why the comet was seen over a rather limited period of time. The striking features of the table are the enormous distances from the Sun, and even more so from the Earth, reached by the submicron-sized dust particles (subjected to a radiation pressure acceleration of 0.6 the solar gravitation acceleration) ejected in the first hours after perihelion. (Such grains dominated the tails of the great Kreutz sungrazers, such as Ikeya-Seki, early after perihelion.) For example, particles ejected a little less than one hour post-perihelion, which in Figure 7 make up the part of the tail appearing just barely above the horizon, were located more than 6 AU from the Sun and more than 7 AU from the Earth on September 2 UT and between 11 and 12 AU on September 28 UT, i.e., the morning hours of September 29 local time, when the comet went out of sight (Ho 1962). On the other hand, the heliocentric and geocentric distances of the microscopic dust ejected as late as 24 hours after perihelion, which occupies the area of the tail near its highest point in Figure 7, were only between 2 and 2.7 AU on September 2 and between 3.6 and 3.9 AU on September 28; of course, the production rate of dust was much higher one hour after perihelion than 23 hours later.
@@
In this context, comparison with the sungrazing comet Pereyra (C/1963 R1) is illuminating. It passed perihelion on September 23, merely three weeks later in the year than predicted for the comet of 1138. Accordingly, Pereyra’s observing conditions were more favorable, but not by much. Also, Pereyra was not a member of Population II, so that the geometry was slightly different. Yet, while intrinsically brighter than Ikeya-Seki, [^6] From sets of visual brightness observations I determined the post-perihelion absolute magnitudes of 5.9 for Ikeya-Seki and 4.4 for Pereyra (Sekanina 2002); these compare with an adopted absolute magnitude of 0.5 for the comet of 1138. Pereyra was not detected in daylight, [^7] If the arrival of comet Pereyra were delayed by 16 years (less than 2 percent of its orbital period), it would have saturated the imager on board the Solwind satellite. but was discovered three weeks after perihelion. Over a period of several days following its discovery, Pereyra was a naked-eye object, displaying a tail several degrees in length. Although the conditions were clearly more favorable in the southern hemisphere (e.g., Bennett et al. 1964), the comet and its tail were seen by northern-hemisphere observers as well. [^8] The visibility of comet Pereyra with the naked eye is apparent from the narratives by some observers [e.g., Capen (1964), who was stationed at Wrightwood, California], but also from an incredible story involving a 1977 letter written by G. E. D. Alcock, in which — as aptly recounted in 2003 by B. G. Marsden in his lecture given in Alcock’s memory to an RAS-BAA Pro-Am Discussion Meeting: Meteorites, Meteors, & Comets, documented by Shanklin (2004) — he wondered, in retrospect, whether a thin pencil-like beam of light that he had seen low in the sky on 1963 September 12, two days before comet Pereyra was discovered, happened to be the comet’s tail. Telescopically the comet was under observation over a period of three months, with the last observation in mid-December made from the northern hemisphere. Thus, notwithstanding the various differences (such as orbital incompatibility, a gap of three weeks in the perihelion time, an intrinsic-brightness gap of almost 4 magnitudes, etc.), one notices very obvious similarities between the predicted traits of the 1138 comet and the observed traits of comet Pereyra.
@@
The second conclusion concerns a rule of thumb to investigate a correlation between the time of last naked-eye observation and the comet’s apparent magnitude. Sightings of a bright sungrazing comet in the ancient or medieval times were those of its tail, which was seen even after the comet’s head was no longer visible to the unaided eye. I tested this relationship on the well-observed sungrazers, such as the Great March Comet of 1843, the Great September Comet of 1882, Ikeya-Seki, etc., and suggested that, based on the experience with these objects, the post-perihelion tail of a sungrazer was typically seen with the naked eye until the brightness of the head diminished to apparent magnitude of approximately 7, thus becoming a telescopic object long before the final tail sighting (Sekanina 2022).
@@
Applied to seven objects between AD 943 and 1695, listed by Hasegawa & Nakano (2001) as potential Kreutz sungrazers (all reaching perihelion under favorable observing conditions in October or November, similar to those for comet Ikeya-Seki in 1965), this exercise led to a revision of their absolute brightness downward, the correction averaging more than 2 magnitudes. The outcome was their better agreement with the absolute magnitudes of the recent Kreutz sungrazers.
@@
A very different case on that same list is the comet of 1041, which was already briefly commented on in Section 2. Hasegawa & Nakano assigned it an absolute magnitude of --1, making it 2 mags intrinsically brighter than was my estimate for the Kreutz system’s progenitor (Sekanina 2022)! There appear to be two problems with this 1041 object, the more serious one being that two different comets — either one recorded in China as well as Korea — were almost certainly involved, as suggested by both Ho (1962) and England (2002). The first of the two comets was discovered on September 1 (Hasegawa & Nakano 2001) and “went out of sight after more than 20 days” (Ho 1962). The mentioned rule-of-thumb method offered for it an absolute magnitude of 4, which was 5 mags fainter than proposed by Hasegawa & Nakano. The second comet appeared in November, was seen for more than 10 days, and appears to have been intrinsically fainter. Both may have been Kreutz sungrazers, perhaps even fragments of an object that split not long before 1041.
@@
However, the estimated (post-perihelion) absolute magnitude 4 for the September comet of 1041 happens to be brighter than Pereyra’s by only 0.5 mag or less and too faint for the parent, especially because circumstantial evidence suggests that the 1041 comet fragmented heavily at perihelion. [^9] The barycentric orbital period of the 1041 comet upon its arrival at perihelion was 678 years, but Pereyra’s original orbital period slightly exceeded 900 years, which implied the existence of one or more fragments in an orbital-period range of $\sim$450–750 years. This would reduce the comet’s rate of post-perihelion fading, thus bringing the relationship between the absolute magnitudes of the 1041 comet and Pereyra closer to expectation. If so, its brightness should have been almost comparable to the brightness of the 1106 or 1138 comets. To achieve this, the time of the last naked-eye observation should have coincided with the comet’s apparent magnitude having equaled not 7, but about 4.5, close to Hasegawa & Nakano’s value. And this is the second problem with the September comet of 1041.
@@ 8. Fragmented Nucleus of the Comet of 1138
Even though I believe that the nucleus of the Great Comet of 1106 did break up [^10] It is possible that besides the Great March Comet of 1843, at least one other sungrazer — C/1668 E1 — was also a fragment of the 1106 comet (Sekanina 2025). in the course of its perihelion passage, only the fragmentation event experienced by the Chinese comet of 1138 is modeled and examined here in some detail. There have been two good reasons: (i) it is virtually certain that the Great September Comet of 1882 and Ikeya-Seki were products of this event (cf. Marsden 1967), possibly with other objects; and (ii) the nuclei of the 1882 comet and Ikeya-Seki were themselves subject to fragmentation at perihelion, the first of the two splitting into five to six components; a search for similarities has been desirable.
The multiple nuclei of the 1882 comet, as well as the existence of split comets whose breakup is known to be of tidal nature but occurred at close proximity of Jupiter — such as D/1993 F2 (Shoemaker-Levy) or 16P/Brooks (C/1889 N1) — suggest that (a) the formation of a chain of more than two nearly equidistant fragments, buried in an oval cloud of debris, is by no means exceptional; and (b) as a rule, the principal fragment is typically located near the chain’s middle, rather than at the leading end, as is usual for nontidally split comets. The conditions (a) and (b) combine to become helpful for predictive purposes. In applications, the first condition serves as an impetus in a search for conceivable fragments among potential historical sungrazers, whereas — once the principal fragment has been identified (or probably identified) — the second condition is useful in assessing a timeline of the fragments in the chain, with the aim to constrain the range of their orbital periods as much as possible. Of particular interest is any hint of evidence that additional fragments could be expected to arrive in the future, especially near future.
@@
In the context of the conditions (a) and (b) above, the adherence to these postulates led to the conclusion that at close proxinity to perihelion the Sun’s tidal force did, in all probability, split the nucleus of the 1138 comet not into two, but at least four or perhaps five or six components (Sekanina 2025). Chronologically, such a chain of fragments could have started with X/1702 D1 and continued with the “sun-comet” of 1792 (Strom 2002), [^12] Because of the very vague nature of the “sun-comets,” this obviously is a weak entry; if it was not a Kreutz member, a fragment expected to arrive at about this time may have been missed. followed by the 1882 sungrazer (the original mass) and by Ikeya-Seki. In order that the principal fragment be located near the middle of the chain, in line with condition (b), one can argue that at least one additional fragment should exist beyond Ikeya-Seki, and it should arrive in the near future. There is a good chance that this could happen — with the uncertainty of at least several years — around 2027, while the next one (if any) should be a matter of distant future.
The appearance of the nucleus region of a split sungrazer is essentially determined by the angular separations between the major, kilometer-sized and larger fragments, which are buried in an encompassing cloud of debris made up of smaller fragments down to microscopic dust. To describe the feature’s overall dimensions, one needs to establish the osculating orbits of the major fragments. Since a tidal breakup that involves no significant momentum exchange among the major fragments has practically no effect on their angular elements, they remain equal to those of the parent at the time of breakup, as does the perihelion time. For the 1138 comet they are listed in Table 1. The two remaining elements vary from fragment to fragment, but the perihelion distances marginally, by a fraction of the nucleus size. The only important differences are in the fragments’ eccentricities, $e$, or alternatively in their semimajor axes, $a$.
The starting point in the process of modeling the appearance of the 1138 comet’s fragmented nucleus is the relationship between its osculating and future orbital periods. From Table 1 the relevant osculating period is 754.4 yr, while taking the 1138 comet as the previous appearance of the 1882 sungrazer, the future barycentric orbital period is 1882.7 – 1138.6 = 744.1 yr. [^13] The actual calculation of the period has to be done in Julian dates (not in calendar days), which are then converted to years. This is necessary because the perihelion times in 1138 and 1882 are given, respectively, in the Julian and Gregorian calendars. The difference does not show up, when a result is rounded off to 0.1 year. In terms of $1/a$, the effect of the planetary perturbations of the 1882 sungrazer’s motion is
@@
Figure 8.— Modeled appearance of the fragmented nucleus of the 1138 comet on 1138 September 2.85 UT, nearly 33 days after perihelion. Plotted are the astrometric positions (equinox J2000) of the proposed major fragments, besides the Great September Comet of 1882 and Ikeya-Seki also the possible fragments X/1702 D1 and the 1792 sun-comet; plus the predicted sungrazer of 2027. The extent of the chain of fragments amounts to ${}^{\prime\prime}\!$.6 in position angle of ${}^{\circ}\!$.3, 77,200 km in space, and is immersed in an elongated cloud of dust debris. All five fragments are assumed to be active at the time to a degree that their condensations are not overwhelmed by the diffused light from the debris, an ideal case discussed below.
@@
The last remaining step is to compute the astrometric positions of the five fragments for the time of interest. Displayed in Figure 8 is the modeled appearance of the nucleus region of the 1138 comet on September 3 around 4:30 local time (September 2.85 UT). The extent of the fragmented nucleus (from X/1702 D1 to the predicted comet of 2027) amounts to ${}^{\prime\prime}\!$.6 in position angle of ${}^{\circ}\!$.3. Terminating the chain of fragments instead with Ikeya-Seki, the length would be ${}^{\prime\prime}\!$.2, with Ikeya-Seki and the 1882 sungrazer only ${}^{\prime\prime}\!$.8 apart. At the geocentric distance of 1.76 AU this was equivalent to merely 3600 km, but because the line of sight made with the line connecting the two objects an angle of merely 14, they were in space 14,500 km apart at the time.
It is illuminating to compare the modeled appearance of the 1138 comet in Figure 8, at a heliocentric distance of 1.12 AU, with the observed appearance of the fragmented nucleus of the 1882 sungrazer at similar post-perihelion distances. Since it will never be known how the nucleus of the 1138 comet actually looked like, this exercise is purely academic. Yet, inspection of the range of possible appearances offers an insight into the potential subsequent evolution and ultimate fate of the fragments.
@@
The first of these reproductions, in Figure 9, shows the fragmented nucleus on 1882 October 13 UT, 25.4 days after perihelion, as seen by Elgin with a 15-cm Grubb equatoreal. The extent of about 20′′ exceeds by 5′′ the length of the modeled chain of fragments of the 1138 comet. As Table 8 explains, this apparent imbalance happens to be the effect partly of a larger geocentric distance of the 1138 object, partly of its highly unfavorable projection, as the chain of fragments made an angle of only 14 with the line of sight. In reality, the 1882 fragmented nucleus was 38,000 km long at the time, less than one half the length of the modeled 1138 nucleus.